OpenStudy (anonymous):

Six friends together earn $167.88 selling crafts. How much money will each friend get if they share the money equally? Complete the equation to model the problem. Use m to represent the unknown value. = 6

OpenStudy (igreen):

All you have to do is divide, so what do you think we could use as an equation?

OpenStudy (igreen):

\(m =\dfrac{?}{?}\)

OpenStudy (anonymous):

divide what

OpenStudy (anonymous):

what do i divide

OpenStudy (anonymous):

@itsbribro

OpenStudy (anonymous):

theres 6 friends total, if they share equally that means they each get the same amount of money so you know you have to divide. the total amount of money goes to each of them equally, so divide 167.88 / 6

OpenStudy (igreen):

$167.88, divided evenly between 6 people, that means you divide 167.88 / 6. Therefore: \(m = \dfrac{167.88}{6}\)